# Beyond Patches: Mining Interpretable Part-Prototypes for Explainable AI

### v1

@@
Abstract.
Deep learning has provided considerable advancements for multimedia systems, yet the interpretability of deep models remains a challenge. State-of-the-art post-hoc explainability methods, such as GradCAM, provide visual interpretation based on heatmaps but lack conceptual clarity. Prototype-based approaches, like ProtoPNet and PIPNet, offer a more structured explanation but rely on fixed patches, limiting their robustness and semantic consistency. To address these limitations, a part-prototypical concept mining network (PCMNet) is proposed that dynamically learns interpretable prototypes from meaningful regions. PCMNet clusters prototypes into concept groups, creating semantically grounded explanations without requiring additional annotations. Through a joint process of unsupervised part discovery and concept activation vector extraction, PCMNet effectively captures discriminative concepts and makes interpretable classification decisions. Our extensive experiments comparing PCMNet against state-of-the-art methods on multiple datasets show that it can provide a high level of interpretability, stability, and robustness under clean and occluded scenarios.
xAI, Model Interpretability, Unsupervised Concept Mining
ccs: Computing methodologies  Knowledge representation and reasoning
1. Introduction
Deep learning (DL) has revolutionized computer vision, enabling models to achieve remarkable performance across a range of tasks, from object detection to image classification. Despite these successes, the opaque nature of DL models has raised concerns regarding their interpretability and trustworthiness. In high-stakes domains such as healthcare (Fellous et al., 2019), autonomous driving (Kim and Joe, 2022), and security (Chen et al., 2019; Nauta et al., 2023), understanding why a model makes a certain decision is as critical as the decision itself. Interpretability not only fosters trust but also aids in identifying potential biases, debugging models, and ensuring compliance with ethical and legal standards. Consequently, there is a growing demand for deep vision models that are not only accurate but also capable of explaining their predictions in a clear and comprehensible manner (Saranya and Subhashini, 2023; Vilone and Longo, 2020).
Common post-hoc explainability techniques such as GradCAM (Chattopadhay et al., 2018; Selvaraju et al., 2017a) or ScoreCAM(Wang et al., 2020) try to interpret the model’s decision by reverse-engineering and finding the regions of the input image that contribute most to the output after the training process, as shown in Fig. 2(a). In contrast to such black-box reverse-engineering methods, ante-hoc techniques (Nauta et al., 2023; Chen et al., 2019; Rymarczyk et al., 2021; Nauta et al., 2021; Zhu et al., 2024) make the backbone produce interpretable components to explain its decision. This recognition-by-components theory (Biederman, 1987) describes how humans recognize objects by segmenting them into multiple meaningful concepts. Some ante-hoc methods such as ProtoPNet (Chen et al., 2019) and PIPNet (Nauta et al., 2023), as shown in Fig. 2(b), segment the spatial deep features extracted by the backbone (before the final pooling layer) into fixed, small patch components and learn a set of prototypes from them. These prototypes, activated by the patches, provide an interpretation of the model’s decision and allow users to identify and compare similar prototypes across different input images. However, these methods have significant limitations. Fixed patch sizes can lead to instability in prototype activation for larger areas, which may fail to focus on primitive semantics. Smaller patches, similarly, may also lose their interpretive capacity due to the limited area they represent. Furthermore, many discriminative and explainable attributes require larger regions to capture their semantic meaning effectively to avoid duplicating the same concept in explanation examples (two similar rows at the bottom of Fig.2(b)). In other words, the regions containing the prototypes can be dynamic and require training.
(a) Heatmap-based
(b) Patched-based
@@ 1. Introduction
Figure 2. Explaining the decision through examples. (a) Heatmap-based methods see the feature backbones as black-box models, and try to locate regions activated by the most important features. (b) With patch-based methods, the features are decomposed into components extracted by image patches, and an explanation is provided by these components.
Direct application of naïve unsupervised part-discovery or Slot Attention-based methods often falls short in capturing primitive, semantically meaningful concepts from images and fails to provide intuitive explanations of model decisions based on extracted prototypes (van der Klis et al., 2023; Locatello et al., 2020). This limitation arises because part features are typically employed merely to identify similar parts across objects, without enforcing conceptual consistency, while Slot Attention features tend to lack explicit prototypical representations required for interpretable reasoning. The result, therefore, is a reduction in the diversity and a lack of interpretability of the prototypes, hindering the model’s ability to provide meaningful and comprehensive explanations for its decisions.
To address these limitations, this paper introduces an intrinsically interpretable image classifier that mines meaningful and interpretable prototypical concepts from dynamic regions in images to explain the model’s decisions. Our unsupervised concept mining network identifies semantically meaningful components, relying only on class labels without requiring additional part labels or annotations. Our part-prototype concept mining network (PCMNet) detects larger, coherent regions of pixels as meaningful prototype parts. This design allows the components to identify relationships between semantically similar prototypes across instances from different classes. Then, we leverage these prototypes to mine a set of primitive yet discriminative concepts used to explain the model’s reasoning directly. For instance, Figure 2 demonstrates an use case where the prototype heatmaps, derived from our pipeline, shows that the decision-making was based on the ”lofty fenders”, ”shiny metal grille” and ”square shaped lights”, to conclude that the vehicle in the image, is a Dodge Durango.
@@
In the first stage, PCMNet employs an unsupervised part discovery mechanism to learn a set of prototypes using a novel center-clustering loss. In the second stage, the model encodes the deep features into a sparse Concept Activation Vector (CAV) (Kim et al., 2018), which provides a spatially and semantically interpretable explanation of the final decision. Each concept activation is computed as the cosine similarity between part features and class-aware prototype centroids. To ensure that these activated concepts are both primitive and class-discriminative, PCMNet applies a clustering algorithm to the part features corresponding to each class-specific prototype. The center of each resulting cluster then serves as a class-aware prototype centroid. Interpretation of the resulting CAV is straightforward, as it is produced by a sparse, non-negative linear layer that guarantees all concept values remain non-negative.
The contributions of this paper are summarized as follows. (1) PCMNet is presented as a part-prototypical concept mining framework for explaining vision classifiers with primitive, prototypical, and interpretable human-meaningful components. (2) Concepts extracted by PCMNet are leveraged from a broader region than single patches to contain more semantical part-prototype abstraction, similar to the human brain, by using a novel contrastive prototype extraction. (3) Two-level clustering mechanisms are proposed for finding interpretable parts at the pixel level and primitive and intuitive concepts at the feature level. (4) Our experiments show that PCMNet explanation is more robust to occlusions, outperforming state-of-the-art patched-based prototype methods according to classification and explainability metrics.
2. Related Work
@@
(a) Post-hoc or heatmap-based explanation: Heatmap-based explainability methods are a widely used family of post-hoc techniques that visualize which parts of an input image contribute most to a model’s decision. These techniques, often referred to as attribution methods, assign importance scores to different image regions to highlight influential features. Gradient-based methods, such as GradCAM (Selvaraju et al., 2017b) and ScoreCAM (Wang et al., 2020), generate heatmaps by backpropagating gradients concerning input features, revealing class-specific activation regions. While GradCAM produces class-dependent heatmaps, other methods like FullGrad (Srinivas and Fleuret, 2019) are class-agnostic, providing a more general view of model behavior across different outputs. Despite their popularity, gradient-based methods suffer from high sensitivity to noise, which can lead to unreliable and inconsistent heatmaps. To mitigate this, gradient-free CAM techniques, such as ScoreCAM, have been proposed to generate more stable and interpretable explanation maps.
@@
Beyond gradient-based methods, attribution propagation approaches offer an alternative way to decompose model predictions into layer-wise relevance scores. Techniques such as Layer-wise Relevance Propagation (LRP) (Bach et al., 2015) recursively distribute relevance through the network, providing a structured breakdown of model decisions. While initially designed for convolutional neural networks (CNNs), recent adaptations have extended these methods to vision transformers (ViTs) (Chefer et al., 2021), leveraging their self-attention mechanisms for improved interpretability. However, despite their effectiveness in highlighting important image regions, both gradient-based and attribution propagation methods remain fundamentally limited in providing high-level, human-interpretable concepts. Unlike our proposed PCMNet, which explains decisions through part-based prototypes, heatmap-based methods do not inherently capture semantic relationships between features, making them less suitable for interpretable concept-based explanations.
Additionally, since these methods are inherently post-hoc, they do not enable real-time user intervention during test-time inference. This means that once a model makes a prediction, users cannot modify or refine the explanations interactively, limiting their practical use in applications requiring human-in-the-loop decision-making. In contrast, PCMNet introduces an ante-hoc, concept-driven interpretability approach that allows for more direct, structured, and interactive explanations by providing semantically meaningful components rather than relying solely on heatmap visualizations.
(b) Slot attention: Slot Attention has emerged as a powerful mechanism for learning object-centric representations by dynamically assigning feature regions to a fixed number of latent slots (Locatello et al., 2020). Unlike traditional attention mechanisms that distribute importance weights across all features, Slot Attention iteratively refines a small set of slots, each representing a distinct object or meaningful component of an image. This allows models to decompose scenes into structured representations without explicit supervision (Kipf et al., 2022). By leveraging iterative updates with attention-based routing, Slot Attention enables more interpretable feature learning, where each slot encapsulates a semantically meaningful part of the input. This property makes it particularly useful for explainable AI, as it provides a structured decomposition of an image, rather than relying on uninterpretable distributed feature maps (Greff et al., 2020).
Recent works have applied Slot Attention in various domains, such as object discovery, segmentation, and interpretable classification (Singh et al., 2022; Monet et al., 2021). In the context of explainable deep learning, it has been used to learn disentangled representations where each slot corresponds to a distinct semantic concept (Karazija et al., 2021). However, these methods typically focus on object-level decomposition rather than part-based representation. Our PCMNet builds on this idea by using Slot Attention-inspired mechanisms to discover meaningful part prototypes rather than full objects, ensuring better granularity in concept-based explanations. By applying slot-like iterative attention within a contrastive prototype extraction framework, PCMNet refines part-based representations, making them more interpretable and aligned with human perception and reasoning.
@@
(c) Concept-based explanation: Contrary to traditional local feature attribution methods that assess the importance of input-level features, concept-based XAI methods (Dreyer et al., 2024; Chen et al., 2019; Rymarczyk et al., 2021; Bach et al., 2015) analyze the role of latent representations in specific layers of a deep neural network (DNN). These methods explore concepts through various mechanisms, including individual neurons, directional representations such as Concept Activation Vectors (CAVs) (Kim et al., 2018), or broader feature subspaces. Early XAI research primarily focused on understanding how these concepts contribute to global decision-making, identifying the most relevant concepts for a given output class (Dreyer et al., 2024; Bach et al., 2015). Some approaches enhance model interpretability by employing local feature attribution techniques to localize and quantify the importance of concepts for individual predictions, thereby enabling concept-based explanations at the instance level (Nauta et al., 2023, 2021). While these methods provide faithful explanations aligned with the model’s internal reasoning, they are vulnerable to performance degradation under occlusion, as their extracted concepts often rely on limited and highly localized discriminative regions of the input. PCMNet addresses this limitation by part-based prototypical concept reasoning to rely on other discriminative parts that are not occluded.
Additionally, although instance-wise concept-based explanations provide deeper insights, they can also be overwhelming and difficult for stakeholders to process, as hundreds of concepts may need to be analyzed for each instance. To address this, some works highlight the benefits of visualizing local decisions within a global embedding space (Dreyer et al., 2024; Donnelly et al., 2022). With PCMNet, we build upon this idea by introducing part prototypes that encapsulate local concepts. This approach reduces the cognitive load for human interpretation by enabling comparisons between individual (local) prediction and prototypical concepts, making explanations more intuitive and structured.
3. Proposed Prototypical Concept Mining Network
@@
Figure 3. Overall architecture of PCMNet. In the first stage, part prototypes are learned from spatial features extracted by the backbone. In the second stage, prototypes are clustered within each class to identify frequently occurring concepts. The center of these clusters is computed as concept prototypes, and the distance between part prototypes and concept clusters determines the activated concepts for the model’s decision-making. This approach enables interpretable and concept-driven classification.
To explain the decision for a classification problem with a training set $\mathcal{T}$ containing $N$ images $\{(x^{1},y^{1}),...,(x^{N},y^{N}\}\in\mathcal{X}\times\mathcal{Y}$, with $L$ classes, the PCMNet extracts primitive and sparse concepts that are activated and discriminate the object. These concepts are described in a limited set of prototypes that existed for almost all images in the datasets (e.g., lights, fenders, and grilles for cars or heads and wings for birds). In other words, the output of our model is the activated concept vector $z_{i}=[z_{p}^{i}]_{p=1}^{K}$, in which $p$ is the prototype index. Fig. 3 shows the overall overview of our approach.
@@
PCMNet transforms features extracted from the backbone model into explainable and interpretable concepts in three steps. Step 1: The model finds semantically independent, meaningful regions from input images in an unsupervised way by minimizing our novel Marginal Cluster Center loss to learn the prototypes that describe these parts. Then, the part features are disentangled to discriminate the class label. Step 2: Once the part features are learned for the training data, for each part inside all instances of each class, a set of shared elementary characteristics is found by applying a clustering approach to these features. Then, the centroid list, which represents the prototypical concepts, is generated by calculating the center of each cluster. Step 3: Extract the Concept Activation Vector (CAV) for the generated centroid list by computing the similarity of extracted features and each centroid. The CAVs are used to discriminate the class of the input and explain the decision. Also, to avoid losing some discriminative concepts that are not prototypical (i.e., specific and unique for this class and not available for others), we combine the features that are located outside of prototypical maps into the CAVs for classification.
3.1. Step 1: Part-Prototype Discovery
@@
Our model is shown in Fig. 3. The input images are given to a backbone to extract the 3D deep features as $F^{i}\in\mathbb{R}^{D\times w\times h}$ where $D$ is the channel size and $w,h$ denotes the width and height. Based on extracted deep features $F^{i}$, our model finds semantically independent meaningful regions that can be shared in different inputs. We name them “part-prototype”. For determining regions, the model scores each pixel in $F^{i}$ to specify their belongingness to each part-prototype class. The mapping score for each pixel in the location $(q)$ is represented by $M^{i}$, we have $\sum_{p=1}^{K+1}M^{i}_{p}(q)=1$. The spatial feature for each part, $F^{i}_{p}$, is computed by element-wise multiplication of $M^{i}_{p}$ to $F^{i}$. The part features, $\mathbf{f}^{i}_{p}$ is computed as:
(1) $\mathbf{f}^{i}_{p}=\mathcal{GP}(\texttt{conv}(F^{i}_{p}))\in\mathbb{R}^{d_{f}},$
@@
where $\mathcal{GP}$ is the global average pooling and $d_{f}$ is the dimension of features. The last index (K+1) is used to cover the background or non-part regions of the foreground (Alehdaghi et al., 2025). We name these features non-prototypical concepts that are computed as:
(2) $\mathbf{g}^{i}=\mathcal{MP}(\texttt{conv}(F^{i}_{K+1}))\in\mathbb{R}^{d_{f}},$
@@
where $\mathcal{MP}$ is the max-pooling. We use the PDiscoNet (van der Klis et al., 2023) as the baseline for extracting parts features. Our part loss comprises their final loss ($\mathcal{L}_{\text{bl}}$) without their orthogonal component and our marginal cluster center loss ($\mathcal{L}_{\text{mcc}}$):
(3) $\mathcal{L}_{\text{part}}=\mathcal{L}_{\text{bl}}+\alpha\mathcal{L}_{\text{mcc}}.$
The goal of $\mathcal{L}_{\text{mcc}}$ is to ensure the selected region for each part remains consistent across all inputs and serves as a prototype; it must capture similar semantic information across different classes. Additionally, these sub-part features should be both class-descriptive and contain unique ID-aware information.
3.1.1. Marginal Cluster Center Loss
We proposed learnable prototypes as features to be distant from each other and then pushed the model to discover the parts from input images to be close to these prototypes. A soft margin is used to make space for the same part from input images with different classes.
(4) $\mathcal{L}_{\text{mcc}}=\sum_{p=1}^{K}[||\mathbf{f}^{i}_{p}-c_{p}||-m_{1}]_{\star}+\frac{1}{K}\sum_{q=p+1}^{K}[m_{2}-||c_{p}-c_{q}||]_{\star},$
where $c_{p}$ is a learnable prototype for part $p$ and $[\cdot]_{\star}$ is $\max(\cdot,0)$.
3.2. Step 2: Part-Centroid List Generation
During training in Step 1, the backbone learns to extract and use part-prototype features from the input data to fulfill its training task. To find all relevant spatial and class-aware concepts, as shown at the top of Fig. 4, we collect all extracted features from the backbone for each class and each part and then cluster them to have a minimal set of plain concepts that consistently describe the visual attributes of those parts. Multiple concepts are considered. They can be seen in each part of each class for more complex objects. To mine those concept prototypes, the centroids of each cluster are computed as $C(l,p,j)\in\mathbb{R}^{d}_{f}$ for a clustering index of $l$, part $p$, and a class of $j$.
3.3. Step 3: Concept Activation Mining

### v2

@@
Abstract
As AI systems grow more capable, it becomes increasingly important that their decisions remain understandable and aligned with human expectations. A key challenge is the limited interpretability of deep models. Post-hoc methods like GradCAM offer heatmaps but provide limited conceptual insight, while prototype-based approaches offer example-based explanations but often rely on rigid region selection and lack semantic consistency. To address these limitations, we propose PCMNet, a part-prototypical concept mining network that learns human-comprehensible prototypes from meaningful image regions without additional supervision. By clustering these prototypes into concept groups and extracting concept activation vectors, PCMNet provides structured, concept-level explanations and enhances robustness to occlusion and challenging conditions, which are both critical for building reliable and aligned AI systems. Experiments across multiple image classification benchmarks show that PCMNet outperforms state-of-the-art methods in interpretability, stability, and robustness. This work contributes to AI alignment by enhancing transparency, controllability, and trustworthiness in AI systems. Our code is available at: https://github.com/alehdaghi/PCMNet.
Introduction
Deep learning (DL) has advanced computer vision, enabling models to achieve remarkable performance across a wide range of tasks, including object detection and image classification (chai2021deep). However, despite these successes, DL models often operate as black boxes, and their lack of transparency has raised critical concerns regarding interpretability and reliability. This problem becomes even more significant in areas such as healthcare (fellous2019explainable), self-driving cars (kim2022xai), and security systems (ProtoPnet; PIPNet), where accurate predictions are not sufficient; it is crucial to understand the reasoning behind them. In these situations, explaining what the model is doing is important. It helps people trust the system, find problems if something goes wrong, and make sure the system follows important rules and laws (markus2021role). Therefore, researchers are now attempting to develop DL models that not only perform well but also provide interpretable and human-understandable explanations for their decisions (saranya2023systematic; vilone2020explainable).
Common post-hoc explainability methods, such as GradCAM (chakraborty2022generalizing) and ScoreCAM (wang2020score), aim to explain a model’s decision by identifying which parts of the input image had the most influence on the output. These methods analyze the trained model’s gradients or activations to identify salient regions of the input. An example is shown in Fig. 1 (a). Unlike these reverse-engineering methods, ante-hoc methods are designed to make the model explainable from the beginning. They guide the model to produce either global (he2025v2c; pcx; jiang2025knowledge) or local (PIPNet; ayoobi2025protoargnet; ProtoTree; zhu2024enhancing) interpretable parts that help explain its decisions. This idea is inspired by the recognition-by-components theory (biederman1987recognition), which suggests that humans understand objects by breaking them down into meaningful parts or concepts.
Some methods, such as ProtoPNet (ProtoPnet) and PIPNet (PIPNet), as illustrated in Fig. 1 (b), divide the spatial deep features (prior to the final pooling layer) into fixed, small patches and learn a set of prototypes from these regions. These prototypes, activated by specific input patches, serve as interpretable visual features that support the model’s decision. This also enables users to compare similar semantic concepts across different images, enhancing transparency and trust in the model’s reasoning capacity. However, these methods have notable limitations. Fixed patch sizes result in unstable prototype activation, particularly for larger semantic regions, where the patches fail to capture coherent or meaningful visual concepts. Conversely, very small patches lack sufficient contextual information, reducing their ability to represent interpretable or semantically rich components. Furthermore, discriminative and interpretable visual concepts often require larger regions to be represented clearly. When the model uses only small, fixed patches, it may fail to capture the full meaning of these concepts. This can lead to repetitive or shallow explanations, as seen in the bottom two rows of Fig. 1 (b), where similar patterns appear multiple times. To address this, the regions used for prototypes should not be fixed. Instead, they should be adjustable and learned during training, allowing the model to focus on more meaningful and diverse parts of the image.
@@ Introduction
Figure 1: Explanation methods. (a) Heatmap-based methods see the feature backbones as black-box models, and try to locate regions activated by the most important features. (b) Patch-based methods decompose the input into image patches and explain the decision using patch-derived components. (c) PCMNet, in contrast, mines both prototypical and non-prototypical concepts, offering a broader and more interpretable set of regions to explain the model decision.
One possible direction is to discover meaningful regions using unsupervised part-based or slot-based mechanisms. However, direct application of naïve part-discovery methods (van2023pdisconet) or Slot Attention techniques (monet2021object; locatello2020slot) fails to produce semantically meaningful and interpretable components. These approaches typically aim to group similar visual parts across instances, but do not enforce conceptual consistency, resulting in ambiguous or fragmented representations (ProtoShare). Moreover, Slot Attention-based features often lack explicit prototypical structures, making it difficult to associate them with clear and intuitive explanations (li2021scouter). As a result, the learned prototypes tend to lack diversity and interpretability, limiting the model ability to generate meaningful justifications for its predictions.
To address these limitations, we propose the Part-Prototypical Concept Mining Network (PCMNet), an interpretable image classifier that explains its predictions by extracting meaningful, part-based prototypical concepts  [^1] We define part-prototypes as localized visual regions that commonly appear across inputs (e.g., a car’s light), while the term concept describes a specific attribute or variation of that part (e.g., rectangular or circular shapes of the light). from selected regions. PCMNet first identifies relationships between semantically related prototypes across different instances, even across class boundaries and then derives a set of primitive class-discriminative concepts to form the basis for the model decision-making process.
@@
In the first stage, PCMNet uses an unsupervised part-discovery module that learns prototypes through a novel center-clustering loss. In the second stage, deep features are encoded into a sparse Concept Activation Vector (CAV) (CAV), which provides a spatially and semantically interpretable representation of the model’s decision. Each concept activation is computed as the cosine similarity between part features and class-aware prototype centroids. To ensure both conceptual simplicity and class-specific relevance, PCMNet applies a clustering algorithm to the part features associated with each prototype, and abstracts each cluster into a representative concept vector. The Concept Activation Vector (CAV) produced by PCMNet is inherently interpretable, as it is generated through a sparse, non-negative linear transformation that ensures all concept scores are positive. As illustrated in Fig. 1 (c), PCMNet activates concepts such as “lofty fenders,” “shiny metal grille,” and “square-shaped lights” to support its prediction that the vehicle is a Dodge Durango, offering a clear and meaningful explanation of the model’s reasoning.
[…]
Related Work
@@
(a) Post-hoc or heatmap-based explanation. Heatmap-based explainability methods are a widely used family of post-hoc techniques that visualize which parts of an input image contribute most to a model’s decision. These techniques, often referred to as attribution methods, assign importance scores to different image regions to highlight influential features. Gradient-based methods, such as GradCAM (selvaraju2017grad) and ScoreCAM (wang2020score), generate heatmaps by backpropagating gradients concerning input features, revealing class-specific activation regions. While GradCAM produces class-dependent heatmaps, other methods like FullGrad (srinivas2019fullgrad) are class-agnostic, providing a more general view of model behavior across different outputs. Despite their popularity, gradient-based methods suffer from high sensitivity to noise, which can lead to unreliable and inconsistent heatmaps. To mitigate this, gradient-free CAM techniques, such as ScoreCAM, have been proposed to generate more stable and interpretable explanation maps.
@@
Beyond gradient-based methods, attribution propagation approaches offer an alternative way to decompose model predictions into layer-wise relevance scores. Techniques such as Layer-wise Relevance Propagation (LRP) (bach2015lrp) or KD-FMV (jiang2025knowledge) recursively distribute relevance through the network, providing a structured breakdown of model decisions. While initially designed for convolutional neural networks (CNNs), recent adaptations have extended these methods to vision transformers (ViTs) (chefer2021transformer), using their self-attention mechanisms for improved interpretability. However, despite their effectiveness in highlighting important image regions, both gradient-based and attribution propagation methods remain fundamentally limited in providing high-level, human interpretable concepts. Unlike our PCMNet, which explains decisions through part-based prototypes, heatmap-based methods do not inherently capture semantic relationships between features, making them less suitable for interpretable concept-based explanations.
@@
(b) Concept-based explanation. Concept-based XAI methods (pcx; ProtoPnet; ProtoShare; bach2015lrp; he2025v2c) analyze the role of latent representations in specific layers of a deep neural network to explore concepts from input images and find Concept Activation Vectors (CAVs) (CAV) for the made decision. Early XAI research primarily focused on understanding how these concepts contribute to global decision-making, identifying the most relevant concepts for a given output class (pcx; bach2015lrp). Some approaches enhance model interpretability by employing local feature attribution techniques to localize and quantify the importance of concepts for individual predictions, thereby enabling concept-based explanations at the instance level (PIPNet; ProtoTree; ayoobi2025protoargnet). While these methods provide faithful explanations aligned with the model’s internal reasoning, they are vulnerable to performance degradation under occlusion, as their extracted concepts often rely on limited and highly localized discriminative regions of the input. PCMNet addresses this limitation by wider range of part-based prototypical concept reasoning to rely on other discriminative parts that are not occluded.
Recently, Concept Bottleneck Models (CBMs) has emerged as a research avenue for inherently interpretable approaches that predict intermediate human-understandable concepts before class prediction (xie2025discovering; posthocCBM; rao2024discover; he2025v2c). Unlike patch- or part-prototype models, CBMs explicitly separate the reasoning process into concept prediction and concept-to-label mapping, allowing direct human intervention on the concept layer. However, their reliance on annotated concepts limits scalability, and the fidelity of the bottleneck depends on concept quality. Recent methods, including post-hoc CBMs (he2025v2c; xie2025discovering) and vision-to-concept tokenizers that learn visual concept vocabularies without text supervision—aim to bridge this gap by enabling scalable and more interpretable reasoning.
Prototypical Concept Mining Network
(a) PCMNet
(b) CMM
@@
Figure 2: (a) Overall architecture of PCMNet. In the first stage, part prototypes are learned from spatial features extracted by the backbone. In the second stage, prototypes are clustered within each class to identify frequently occurring concepts. The center of these clusters is computed as concept prototypes, and the distance between part prototypes and concept clusters determines the activated concepts for the model’s decision-making. This approach enables interpretable and concept-driven classification. (b) Concept Mining Module. First, a part centroid list is generated by applying DBSCAN(DBSCAN) clustering within each class and computing the centers of the resulting clusters. These centroids serve as frequently occurring concept prototypes. Next, to activate the most relevant concept for a given part feature, the similarity (inverse distance) to the centroids is measured. These values are then used as CAVs to inform the model’s decision-making.
To explain classification decisions over a training set $\mathcal{T}={(x^{1},y^{1}),...,(x^{N},y^{N})}\subset\mathcal{X}\times\mathcal{Y}$ with $L$ classes, PCMNet extracts sparse and primitive concepts that are both activated and discriminative for object recognition. These concepts are captured through a compact set of prototypes that recur across images in the dataset (e.g., lights, fenders, and grilles for cars; heads and wings for birds). The model outputs a class-based Concept Activation Vector (CAV) $\mathbf{z}^{i}=[\mathbf{z}^{i}_{1};\dots;\mathbf{z}^{i}_{K}]$, where ; denotes concatenation and each $\mathbf{z}_{p}$ represents the activated concepts from part $p$.
@@
Fig. 2 shows the overall overview of PCMNet, which transforms features extracted from the backbone into explainable and interpretable concepts in three steps. Step 1: PCMNet identifies semantically meaningful and independent regions from input images in an unsupervised manner by minimizing the Marginal Cluster Center loss ($\mathcal{L}_{\text{mcc}}$). This process learns part-prototype features $\mathbf{f}^{i}_{p}$ that are informative for class discrimination. Step 2: Once part-level features are obtained, the Concept Mining Module (CMM) $\mathcal{H}$ aggregates these features across all instances of each class to discover a set of shared elementary visual patterns. Cluster centers derived from this process form the list $\mathcal{C}$, representing the core prototypical concepts. Step 3: For each input, a Concept Activation Vector (CAV) $\mathbf{z}^{i}$ is computed by measuring the similarity between its extracted part features and the concept list $\mathcal{C}$. These CAVs are used both for classification and for explaining the model’s decision. To preserve class-specific but non-shared discriminative cues, PCMNet also incorporates features outside the prototypical regions into the CAVs.
Step 1: Part-Prototype Discovery
@@
Our model is shown in Fig. 2. The input images are given to a backbone to extract the 3D deep features as $F^{i}\in\mathbb{R}^{D\times w\times h}$ where $D$ is the channel size and $w,h$ denote the width and height. Based on extracted deep features $F^{i}$, our model finds semantically independent meaningful regions that can be shared in different inputs. We name them “part-prototype”. For determining regions, the model scores each pixel in $F^{i}$ to specify their belongingness to each part-prototype class. The mapping score for each pixel $(a)$ is represented by $M_{p}^{i}(a)$, we have $\sum_{p=1}^{K+1}M^{i}_{p}(a)=1$. The spatial feature for each part, $F^{i}_{p}$, is computed by element-wise multiplication of $M^{i}_{p}$ to $F^{i}$. Then part features, $\mathbf{f}^{i}_{p}$ is computed as:
$\vskip-2.84544pt\mathbf{f}^{i}_{p}=\mathcal{GP}(\texttt{conv}(F^{i}_{p}))\in\mathbb{R}^{d_{f}},$ (1)
@@
where $\mathcal{GP}$ is the global average pooling and $d_{f}$ is the dimension of features. The last index (K+1) is used to cover the background or non-part regions of the foreground (BMDG). We name these features non-prototypical concepts that are computed as:
$\vskip-5.69046pt\mathbf{g}^{i}=\mathcal{MP}(\texttt{conv}(F^{i}_{K+1}))\in\mathbb{R}^{d_{f}},$ (2)
@@
where $\mathcal{MP}$ is the max-pooling. We use the PDiscoNet (van2023pdisconet) model for extracting parts features. Our part loss comprises their final part-discovery loss ($\mathcal{L}_{\text{pd}}$) without their orthogonal component and with our marginal cluster center loss ($\mathcal{L}_{\text{mcc}}$):
$\vskip-4.26773pt\mathcal{L}_{\text{part}}=\mathcal{L}_{\text{bl}}+\alpha\mathcal{L}_{\text{mcc}}.$ (3)
where $\alpha$ is a parameter regularizing the effect of $\mathcal{L}_{\text{mcc}}$. The goal of $\mathcal{L}_{\text{mcc}}$ is to ensure that each selected region is consistent across inputs and captures semantically similar content, while remaining class-discriminative and identity-aware.
Marginal Cluster Center Loss
To ensure part-prototypes are both distinct and semantically meaningful, we encourage separation in the feature space while guiding part-level features to align with their corresponding prototypes. To handle intra-part variation across classes, we apply a soft-margin loss that allows features from the same part (but different classes) to stay within a bounded distance of their prototype. This encourages the model to discover consistent, class-aware part structures while preserving inter-class separation between prototypes. The loss can be written as
$\mathcal{L}_{\text{mcc}}\!=\!\sum_{p=1}^{K}[|\mathbf{f}^{i}_{p}-c_{p}|\!-\!m_{1}]_{\star}\!+\!\frac{1}{K}\!\sum_{q\neq p}^{K}[m_{2}\!-\!|c_{p}-c_{q}|]_{\star},$ (4)
where $c_{p}$ is a learnable prototype and $\cdot$$_{\star}$ is $\max(\cdot,0)$.
Step 2: Part-Centroid List Generation
In Step 1, the backbone learns to extract meaningful part-level features that support the classification task. To discover compact and consistent class-specific concepts, we collect all part features from the training set. Specifically, for each class $j$ and part $p$, we construct $B^{j}_{p}=\{\mathbf{f}_{p}^{i}\mid y^{i}=j\}\in\mathbb{R}^{N_{j}\times d_{f}}$, where $N_{j}$ is the number of training samples from class $j$. To identify semantically coherent patterns, we apply the clustering algorithm separately to each $B^{j}_{p}$, yielding $T^{j}_{p}\in\mathbb{N}^{N_{j}}$, which assigns cluster labels to the part features. We then compute the centroid of each cluster $l$ (representing a visual concept) as:
$\vskip-2.84544pt\mathcal{C}(l,p,j)=\text{mean}(\{B^{j}_{p}[T^{j}_{p}=l]\})\in\mathbb{R}^{d_{f}}.$ (5)
The total number of discovered concepts across all classes and parts is computed as $d_{c}=\sum_{j=1}^{L}\sum_{p=1}^{K}\max(T^{j}_{p})$. This process, shown in the top of Fig. 2(b), allows the model to build a compact and expressive vocabulary of class-specific, part-aware visual concepts.
Step 3: Concept Activation Mining